Circular No. 7721

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001dj IN NGC 180 A. V. Filippenko and R. Chornock, University of California at Berkeley,

report: “Inspection of CCD spectra (range 330–1000 nm) obtained on Sept. 20 UT with the Shane 3-m reflector at Lick Observatory reveals that SN 2001dj (IAUC 7674) is indeed a supernova, contrary to the report on IAUC 7680. The bright star located 56′′.6 east and 3′′.8 north of the supernova has an early-F spectral type, so we suspect this is the star whose spectrum is described on IAUC 7680. The supernova is a peculiar type-II event in the nebular phase, with strong, broad Hα emission having an asymmetric profile (the line peaks at −5000 km/s, after removal of the NED redshift); it also shows reasonably strong but somewhat narrower He I emission lines. He I emission lines at this phase are very rare in type-II supernovae.”

SUPERNOVAE 2001di, 2001dv, 2001ei, AND 2001ej Filippenko and Chornock add: “Inspection of CCD spectra, obtained

as above on Sept. 20 UT, reveals that SN 2001dv (IAUC 7690) is of type II, perhaps a month after maximum brightness; Hα emission dominates the spectrum. SN 2001ei (IAUC 7718) is a peculiar type-Ia event, ∼ 3 weeks past maximum brightness, spectroscopically very similar to the highly subluminous SN 1991bg (Filippenko et al. 1992, A.J. 104, 1543); it even exhibits the narrow Na I D absorption that was seen in SN 1991bg but not seen in most other subluminous type-Ia supernovae. SN 2001di (IAUC 7673) is a type-II supernova beginning to enter the nebular phase; Hα emission is very strong. SN 2001ej (IAUC 7719) is a very young type-Ib or -Ic supernova with a blue continuum and relatively shallow absorption lines; further monitoring should facilitate a definitive classification based on the presence or absence of He I lines.”

T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001ej, obtained by M. Calkins on Sept. 20.50 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ic supernova near maximum light.

SUPERNOVA 2001el IN NGC 1448 Libert A. G. Monard (not Liebert, as given on IAUC 7720) reports

that SN 2001el has brightened to m_{v} = 13.7 on Sept. 20.114 UT.

2001 September 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green